OpenStudy (anonymous):

4e*t x e*t=?? * is exponent

OpenStudy (jdoe0001):

\(\bf 4e^t\times e^t\quad \textit{recall}\quad a^n\times a^m=a^{n+m}\\ \quad \\ 4e^t\times e^t\implies 4e^{t+t}\)